OpenStudy (anonymous):

I have the answer can someone please check me.... Which function has the following as its derivative b'(x)=cos(3x)

OpenStudy (anonymous):

\[ b(x)=\int\cos(3x)dx\\ \text{set}\quad u=3x\implies dx=\frac{du}{3}\\ b(u)=\int\cos(u)\frac{du}{3}=\frac{\sin u}{3}+C \]

OpenStudy (anonymous):

I got (1/3)sin(3x) is that correct

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\sin(3x)\]

OpenStudy (anonymous):

yes... do not forget the "+C" it is indefinite integrals

OpenStudy (anonymous):

what do you mean can you show me what is would look like

OpenStudy (anonymous):

can someone please tell me if the answer I gave is correct

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@MATTW20 you are on the opposite side of the river.

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

@MATTW20 re-read the question @farmergirl411 you are good. what you go is correct.

OpenStudy (anonymous):

oh my fault misread

OpenStudy (anonymous):

so then cansomeone please clarify what the answer is

OpenStudy (anonymous):

you had the right answer

OpenStudy (anonymous):

(1/3)sin(3x) is that what is right

OpenStudy (anonymous):

@MATTW20 do not confuse the poor childrean!! ;ol

OpenStudy (anonymous):

lol yeah sorry bout that

OpenStudy (anonymous):

@farmergirl411 Remember, that for "indefinite integrals", i.e., when the integration has no bounds, you always have to retricean unknown constant term (call it C)

OpenStudy (anonymous):

so is what I posted what the answer is correct when you say the C part I am getting really confused

OpenStudy (anonymous):

so, you exact answer should be.. \[ b(x)=\frac{1}{3}\sin(3x)+C \] VERIFICATION: try to find b'(x).. should give you the question back...

OpenStudy (anonymous):

ok thanks